### v2

@@ Abstract
Generative machine learning models on sequences are transforming protein engineering. However, no principled framework exists for conditioning these models on auxiliary information, such as experimental data, in a plug-and-play manner. Herein, we present ProteinGuide—a principled and general method for conditioning—by unifying a broad class of protein generative models under a single framework. We demonstrate the applicability of ProteinGuide by guiding two protein generative models, ProteinMPNN and ESM3, to generate amino acid and structure token sequences, conditioned on several user-specified properties such as enhanced stability, enzyme classes, and CATH-labeled folds. We also used ProteinGuide with inverse folding models and our own experimental assay to design adenine base editor sequences for high activity.
@@
Generative models are enabling protein engineers to create bespoke proteins for therapeutic, agricultural and environmental uses [1, 2, 3]. The outcome of any AI-based protein engineering approach is to ultimately specify a protein sequence, or a set of such sequences, even when designed by way of structure. For example, one might first use a backbone structure generative model such as RFdiffusion [2], Chroma [3], or ESM3 [4], followed by use of a backbone-conditioned generative model on sequences (also called an inverse-folding model), such as ProteinMPNN [1] or ESM-IF1 [5], to generate sequences likely to fold into the specified backbones. Alternatively, one might forgo generating a backbone structure intermediary and instead generate sequences directly from an “evolutionary” sequence-only model such as ESM3 [4], ProGen [6], EvoDiff [7], DPLM [8] and others [9, 10]. In all cases, the final step employs a generative model of sequences.
A key ingredient missing from the current protein engineering AI menu is the plug-and-play ability to post hoc “guide” sequence generative models toward properties of interest, because the technical underpinnings to do so have remained elusive. By guide, we mean, formally, to condition the probability distribution encoded by an existing generative model, $p(x)$, on newly-desired properties of interest, $p(y\in Y|x)$, so as to be able to generate (i.e., sample) sequences from the desired conditional distribution, $x\sim p(x|y\in Y)$. Importantly, we want to condition in a plug-and-play manner, meaning that we want to avoid re-training the existing generative model, instead modulating how we sample from it with an auxiliary model, such as a property predictive model. Although it is currently possible to re-train generative models to achieve property conditioning [11, 12], the plug-and-play characteristics of guidance would enable everyday users to readily adapt existing large, expensive-to-train models of interest to suit their needs, on-the-fly [13, 14, 15]. [^1] Herein, we use the term “guide” and “guidance” to refer to “predictor guidance” in Nisonoff et al. [16]. Interested readers can also find how to perform predictor-free guidance in that same work.
Figure 1: Equivalence between classes of generative models enables guidance with ProteinGuide. a. Schematic showing how Masked Language Models (e.g., ESM3) and order-agnostic autoregressive models (e.g., ProteinMPNN) can be viewed through the lens of masking diffusion/flow models. This interpretation enables generalization of our guidance method [16] to obtain ProteinGuide. Masking of one sequence position is denoted with the symbol “?”; colored dots help to visually tag individual sequences. The sequence being masked during training is M K R S V T; $p_{\theta}(\mathbf{x})$ denotes the pre-trained generative model. The shared training objective involves randomly masking out parts of the sequence and then predicting the masked sequence given the remaining unmasked context, over different masking rates denoted by more or fewer “?” in a sequence. b. Contrasting unguided generation with guided generation by way of ProteinGuide. A sequence is denoted, $\mathbf{x}$, the pre-trained generative model is denoted $p_{\theta}(\mathbf{x})$; $y$ denotes a property of interest that we seek to guide toward using predictive model $p_{\phi}(y\mid\mathbf{x})$. When guiding, we coherently blend information from the predictive model with that of the pre-trained model in a statistically correct manner. Red letters indicate mutations that are introduced when guiding with the predictive model, and not seen when generating in an unguided manner.

### v4

@@ Abstract
Sequence generative models are transforming protein engineering. However, no principled framework exists for conditioning these models on auxiliary information, such as experimental data, without additional training of a generative model. Herein, we present ProteinGuide, a method for such “on-the-fly” conditioning, amenable to a broad class of protein generative models including Masked Language Models (e.g. ESM3), any-order auto-regressive models (e.g. ProteinMPNN) as well as diffusion and flow matching models (e.g. MultiFlow). ProteinGuide stems from our unifying view of these model classes under a single statistical framework. As proof of principle, we perform several in silico experiments. We first guide pre-trained generative models to design proteins with user-specified properties, such as higher stability or activity. Next, we design for optimizing two desired properties that are in tension with each other. Finally, we apply our method in the wet lab, using ProteinGuide to increase the editing activity of an adenine base editor in vivo with data from only a single pooled library of 2,000 variants. We find that a single round of ProteinGuide achieves a higher editing efficiency than was previously achieved using seven rounds of directed evolution.
@@
Generative models are enabling protein engineers to create bespoke proteins for therapeutic, agricultural and environmental uses [20, 119, 56]. The outcome of any AI-based protein engineering approach is to ultimately specify a protein sequence, or a set of such sequences, even when designed by way of structure. For example, one might first use a backbone structure generative model [119, 56, 44], followed by use of a backbone-conditioned generative model on sequences (also called an inverse-folding model) [20, 53], to generate sequences likely to fold into the specified backbones. Alternatively, one might forgo generating a backbone structure intermediary and instead generate sequences directly from a sequence-only model [44, 76, 1, 118, 77, 93]. In all cases, the final step employs a generative model of sequences.
In most practical settings one is interested in generating sequences with particular properties of interest—often properties that are rare in the distribution of natural sequences. For example, suppose one wishes to increase the catalytic activity of an enzyme for some substrate, then these desired proteins are not expected to be well-represented in most general-purpose, pre-trained generative models. Nor are they likely to be well-represented by conditioning on existing annotation labels in such models. Consequently, protein engineers often acquire experimental measurements to glean which proteins are more likely to express this property at the desired level. With these newly generated data, they can construct a predictive model for the measured property, with the hope of leveraging this model to generate new sequences to better achieve their design goal. However, current strategies for incorporating such a model or other auxiliary information into generation remain limited and often ad hoc.
A common workaround to re-training a generative model is to use post hoc filtering—that is, to generate many candidate sequences from a general, pre-trained model and then to filter these sequences according to a predictive model [128, 6]. However, post hoc filtering is fundamentally constrained by the pre-trained model’s distribution, which makes sampling of desired proteins increasingly inefficient as the desired sequences become more rare in the pre-trained generative model distribution [71]. Moreover, for filtering on more than one property, the sampling inefficiency problem is compounded [48]. Alternatively, fine-tuning—which does not modulate the sampling distribution on-the-fly and does require re-training a generative model—is another commonly used approach [133, 84]. In fine-tuning, a set of proteins thought to most exhibit the desired property—such as those that scored most highly in the laboratory—are used to continue training the pre-trained generative model [92, 81, 57]. Fine-tuning concentrates the generative model distribution on the fine-tuning dataset, thereby coercing it to sample from the relevant part of the existing pre-trained model distribution, rather than enabling extrapolation beyond that distribution [81, 80, 92]. Furthermore, fine-tuning risks overwriting important general domain information learned by the model [64, 75], such as features related to the stability of a protein. The pre-trained model can learn about such general properties from data sets that are typically orders of magnitude larger than the fine-tuning data set. We shall see the practical difficulties that fine-tuning encounters in our experiments. See “Alternative approaches for infusing generative models with auxiliary information” and Section 2.1 for more detailed discussions of other methods.
The aforementioned limitations point to the need for a methodological framework that can infuse sequence generation from a general, pre-trained model with information about the desired properties, all-the-while retaining general knowledge absorbed into the pre-trained model. Development of such a framework should increase our chances of discovering parts of protein design space that are inaccessible to the generative model alone, or to the predictive model alone. We would also like this framework to handle multiple properties, which might be in tension with one another, such as enzyme specificity and activity.
Herein, we propose ProteinGuide (Fig. 1) to fill this gap. Specifically, ProteinGuide enables “on-the-fly” conditioning, such that we do not train a new generative model; instead, we iteratively modulate, or “guide” the generation process of a pre-trained unconditional model with information from one or more property predictive models. Each sampling step in this guidance procedure represents a sampling step from the desired conditional generative model, without the need to re-train a generative model to explicitly capture the desired conditional distribution.